The perimeter of a rectangle is 28cm. The length x, is 8cm more than the width, Y. Write two equations that could be used to solve the system. Solve for Y and Graph.
Answer by stanbon(75887)

A concert hall has 100 seats in the first row, 102 seats in the second row, 104 in the third row, and so on. The last row has 168 seats.
-------
Arithmetic series with a(1) = 100 and d = 2
Find "n"::
a(n) = 2 + (n-1)2 = 168
(n-1)2 = 166
n-1 = 84
n = 85
----------------
a) how many total rows does the concert hall have?
Ans:: 85
--------------------
B) how many seats does the concert hall have?
S(85) = (85/2)(100 + 168) = 11390
-------
Cheers,
Stan H.
-----------